OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <2, -4>, v = <3, -8>

OpenStudy (jdoe0001):

\(\bf \textit{angle between two vectors }\\ \quad \\ cos(\theta)=\cfrac{u \cdot v}{||u||\ ||v||} \implies \cfrac{\text{dot product}}{\text{product of magnitudes}}\\ \quad \\ \theta = cos^{-1}\left(\cfrac{u \cdot v}{||u||\ ||v||}\right)\)

OpenStudy (anonymous):

Is it neither?

OpenStudy (jdoe0001):

\(\huge ?\)

OpenStudy (anonymous):

my bad I meant 6.0 degrees?

OpenStudy (jdoe0001):

hmm what did you get for the dot product?

OpenStudy (anonymous):

-4(-8)

OpenStudy (jdoe0001):

\(\bf <a,{\color{blue}{ b}}>\cdot <c,{\color{blue}{ d}}>\implies (a\cdot c)+({\color{blue}{ b}}\cdot {\color{blue}{ d}})\)

OpenStudy (jdoe0001):

\(\bf <a,b>\qquad magnitude\implies \sqrt{a^2+b^2}\)

OpenStudy (anonymous):

So what would the answer come out to be?

OpenStudy (jdoe0001):

well. divide the dot product by the magnitudes product then get the \(\bf cos^{-1}\)

OpenStudy (anonymous):

3.0 degrees?

OpenStudy (jdoe0001):

hmm dunno

OpenStudy (anonymous):

k, I'm pretty sure my first answer was correct. Thanks for the help

OpenStudy (jdoe0001):

yw